OpenStudy (anonymous):

Triangle DEF is similar to triangle PQR. Side DE measures 3, side EF measures 7 and side FD measures 9. Side PQ measures 5x – 25 and side RP measures 10x – 5. Find the value of x.

OpenStudy (anonymous):

Similar triangles mean that you can set the sides into a fraction and set them equal: \[\frac{5x -25}{3} = \frac{10x - 5}{9}\] solve for x from this